[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Dec 05, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Dec 5 16:49:44 UTC 2016

Here is today's summary of NHC computer operations:
--- Craig Mattocks


No new issues.
No new issues.
Other items:
1. There will be a WCOSS production switch tomorrow, Tuesday, December 6th,  
from Orlando, FL (Gyre/Surge) to Reston, VA (Tide/Luna) for security  
patching (Dirty COW, etc.).

2. NCO network engineers upgraded the NCEP network to 100G last week. Today  
a new WAN router will be installed in Reston and some firewall work will be  
performed. During the week of December 12th, a 30-60 minute disruption may  
occur when equipment is upgraded in Orlando as part of the project. NCO is  
evaluating the new performance levels and is adjusting throttling to  
maximize data throughput.

3. At 21 UTC yesterday, the NCEP Real Time Data Monitoring System (RTDM)  
detected a drop in the amount of high-resolution GOES soundings and  
radiances that were received and assimilated into the GFS (82.6% of normal)  
and NAM (64.5% of normal). The dataflow returned to normal after 00 UTC  
1. RFC #211 - ksh, Gempak/N-AWIPS v1.1.0 - Update TAFB's High Wind and Seas  
graphical product generation Gempak Korn shell scripts to trim/shrink the  
northern extent of the domain. This change, requested by the TAFB Chief,  
will bring the High Wind domain into better alignment with TAFB's AOR (area  
of responsibility). To be implemented on Monday, December 5 at 1500Z (10 am  

2. RFC #2644 - Pull in ASCAT data from the the Domestic Data Service (DDS),  
write it to disk (on the Compute Farm, at the request of the National  
Centers) and send it to NHC systems. To be implemented on Thursday,  
December 8 at 1500Z (10 am EST).

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161205/9415e0bc/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list